OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 30 = x2 + x? A. (−5, 0) and (6, 0) B. (−5, 0) and (−6, 0) C. (5, 0) and (−6, 0) D. (5, 0) and (6, 0)

OpenStudy (anonymous):

x int is where y = 0

OpenStudy (anonymous):

x2 + x - 30 = 0 (x-5) (X+6) = 0 x = 5 and x = -6